OpenStudy (anonymous):

if a degree on a right triangle is 62, its adjacent length is 15ft and its opposite angle is x. what is the length of the opposite side??

OpenStudy (amistre64):

angle(x)= 90-62 degrees.

OpenStudy (amistre64):

Law of sines works: sin(62)/a = sin(90-62)/15 a= 15sin(62)/sin(28)

OpenStudy (amistre64):

a = about 28.21 ft if I did it right :)

OpenStudy (anonymous):

thank you!

OpenStudy (amistre64):

was I right?

OpenStudy (anonymous):

yep

OpenStudy (amistre64):

yay!! I like being right, makes me feel smart...then I lock my keys in the car and feel stupid again lol

OpenStudy (anonymous):

haha wow